I need help with these problems please explain them to me step by step thanks....I have an exam tomorrow i want to really understand them;.....
Problem #1
Find the percent of the total area under a normal curve that is contained within the specified interval. Between the mean and 1.64 standard deviations from the mean
Problem #2
Find the percent of the total area under the standard normal curve between the given z-scores.
z= -1.10 and z = -0.36

Answer by stanbon(48568)

Problem #1
Find the percent of the total area under a normal curve that is contained within the specified interval. Between the mean and 1.64 standard deviations from the
You need a z-chart or a calculator with z capability, like a TI 83 or 84
On the z cart find z=1.64. The chart tells you the area over the interval
[0,1.64] = 0.9495-0.5000 = 0.4495
--------------------
Problem #2
Find the percent of the total area under the standard normal curve between the given z-scores.
z= -1.10 and z = -0.36
Look up z = 1.10 = 0.8643-0.5000 = 0.3643
Look up z = 0.36 = 0.6406-0.5000 = 0.1406
Subtract the 2nd area from the 1st to get 0.2237
--------
Comment: Your z-chart might be arranged differently
but the final figure you get should be the same
as the one I have.
===============
Cheers,
Stan H.